OpenStudy (devonhoward15):

The variable y varies inversely with x and y=8 when x=4. what is y when x=2

OpenStudy (devonhoward15):

@Daniellelovee

OpenStudy (devonhoward15):

i think its 4 cause y is double x and 2+2=4

OpenStudy (devonhoward15):

that how i got my anwser lol... i read that

OpenStudy (daniellelovee):

oh ok good let me check y = a/x 8=a/4 a=2 y=2/x

OpenStudy (daniellelovee):

and now just replace :)

OpenStudy (devonhoward15):

so it is 4

OpenStudy (devonhoward15):

right?

OpenStudy (ijlal):